A rotating wheel has a radius of 6 ft and 6 in. A point on the circumference of the

wheel moves 30ft and 2 seconds. Find the angular velocity of the wheel?


1
Expert's answer
2021-03-14T19:18:57-0400

"\omega=\frac vr,"

"v=\frac lt"

"\omega=\frac{l}{rt}=\frac{30}{6.5\cdot2}=2.3~\frac{rad}{s}."